A student begins with 125 mL of a Na2SO4 solution and evaporates the water from the mixture. The resulting material has a mass of 19.83 g.  What is the concentration (M) of the original solution?

(calculate answers to 2 decimal places)


1
Expert's answer
2021-10-26T04:40:02-0400

molar mass of "Na_2SO_4" = 142 g/mol

mass of "Na_2SO_4" in the solution = 19.83 g

number of moles of "Na_2SO_4" = "\dfrac{19.83}{142}" = 0.14 mol


Volume of solution = 125 mL = 0.125 L


molarity (molar concentration) ="\dfrac{\text{number of moles}}{\text{volume of solution}}" = "\dfrac{0.14}{0.125}" = 1.12 M"."